How does the angle of an inclined plane determine the friction on an object that is on the plane?
 
 

Grade:8

2 Answers

Ojas
5 years ago
Let angle of an inclined plane be x. Then tanx will be equal to the coefficient of friction. 

coefficient of friction is a value that shows the relationship between the force of friction between two objects and the normal reaction between the objects that are involved. It is a value that is sometimes used in physics to find an object's normal force or frictional force when other methods are unavailable.

The coefficient of friction is shown by {\displaystyle F_{f}=\mu F_{n}\,}{\displaystyle F_{f}=\mu F_{n}\,} . In that equation, {\displaystyle F_{f}}{\displaystyle F_{f}} is the frictional force{\displaystyle \mu }\mu  is the coefficient of friction, and {\displaystyle F_{n}\,}F_{n}\, is the normal force.

The coefficient {\displaystyle \mu }\mu can be two different things. It is either the coefficient of static friction {\displaystyle \mu _{s}}\mu_s, or the coefficient of kinetic friction {\displaystyle \mu _{k}}\mu_k.The coefficient of static friction is the friction force between two objects when neither of the objects is moving. The coefficient of kinetic friction is the force between two objects when one object is moving, or if two objects are moving against one another.

Atul Nagar
5 years ago
The coefficient of Friction is given by Tan A, if A is an angle of incline.
Proof : The normal reaction is N = Mg Cos A
Therefore , if the object on the incined is in equilibrium, then component of weight along incline = force of friction
mg Sin A = f , force of friction
f= u N  (u = coeff of friction)
f= u Mg Cos A
 
hence; Mg Sin A = u Mg Cos A
hence Tan A = u (Coefficient of friction.)
 
Askiitians member 
 
  1. Tech Chem